OpenStudy (anonymous):

What is y+1=-4(x-3) rewritten in General Form (Ax + By + C = 0)?

OpenStudy (texaschic101):

y + 1 = -4(x - 3) --- first we distribute through the parenthesis y + 1 = -4x + 12 --- subtract 1 from both sides y = -4x + 12 - 1 -- simplify y = -4x + 11 -- now add 4x to both sides 4x + y = 11 -- now subtract 11 from both sides 4x + y - 11 = 0

OpenStudy (anonymous):

Thank you SO much. I can solve up until distributing through the parenthesis and then I get lost LOL Thank you so much again!!!! xo

OpenStudy (texaschic101):

no problem...it can get quite confusing at times :)